# Causality Is Key to Understand and Balance Multiple Goals in Trustworthy ML and Foundation Models

### v4

@@ 4 Challenges and Opportunities
Challenges in Causal Foundation Models. One foundation model-specific challenge is concept superposition, particularly in LLMs, where multiple meanings are entangled within a single representation, complicating causal reasoning (Elhage et al., 2022). Vision models exhibit this issue to a lesser extent due to their structured data formats. Being aware of superposition is imporant for effectively integrating causality.
@@
Another challenge is the lack of high-quality causal data. Training foundation models with causal reasoning requires datasets annotated with explicit causal structures or interventional data, which are scarce and expensive to produce. Scalable methods for generating synthetic causal datasets show a promising direction. Alternatively, focusing on post-training methods allows causal interventions in a more data-efficient way.
@@
Additionally, the computational complexity of integrating causal reasoning into foundation models poses a significant challenge. For fine-tuning, low-rank adaptation methods such as LoRA can be employed to reduce the number of learnable parameters, making causal integration more efficient without compromising performance (Hu et al., 2021).
@@ 5 Alternative View
Some may argue that different domains prioritize different requirements for trustworthy ML, and there is no need to reconcile them. However, this perspective is unlikely to hold universally, as most real-world applications intersect with multiple ethical and trustworthy ML principles, such as fairness, privacy, and robustness, which must be balanced to ensure reliable outcomes.
Another perspective suggests that causal properties can emerge spontaneously by training larger models on vast amounts of data. While this is possible, it is not guaranteed, and more importantly, it provides no control over whether or how these properties arise. In contrast, much of the scientific causal knowledge already exists, and finding ways to integrate this knowledge with machine learning models offers a more resource-efficient, reliable, and explainable pathway to achieving trustworthy ML.
@@ 6 Conclusion and Call for Action
7 Impact Statement
This paper advocates for integrating causality into foundation models to enhance fairness, privacy, robustness, and explainability. By reducing reliance on spurious correlations and improving decision-making, causal methods can make AI systems more reliable, transparent, and aligned with human values—especially in high-stakes domains like healthcare, law, and finance. Adopting causality-driven AI has the potential to improve trust, regulatory compliance, and ethical governance, ultimately contributing to a more fair, transparent, and socially beneficial technological landscape.
Acknowledgements
This work is funded in part by PriSyn: Representative, synthetic health data with strong privacy guarantees (BMBF), grant No. 16KISAO29K. It is also supported by Integrated Early Warning System for Local Recognition, Prevention, and Control for Epidemic Outbreaks (LOKI / Helmholtz) grant. The work is also partially funded by Medizininformatik-Plattform ”Privatsphären-schutzende Analytik in der Medizin” (PrivateAIM), grant No. 01ZZ2316G, and ELSA – European Lighthouse on Secure and Safe AI funded by the European Union under grant agreement No. 101070617. Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or European Commission. Neither the European Union nor the European Commission can be held responsible for them.

### v5

@@ 4.2 Challenges in Causal Foundation Models
Concept Superposition. One foundation model-specific challenge is concept superposition, particularly in LLMs, where multiple meanings are entangled within a single representation, complicating causal reasoning elhage2022toy . Vision models exhibit this issue to a lesser extent due to their structured data formats. Being aware of superposition is imporant for effectively integrating causality.
@@
Lack of High-quality Causal Data. Training foundation models with causal reasoning requires datasets annotated with explicit causal structures or interventional data, which are scarce and expensive to produce. Scalable methods for generating synthetic causal datasets show a promising direction webster2020measuring ; chen2022disco . Alternatively, focusing on post-training methods allows causal interventions in a more data-efficient way.
@@
Computational Complexity. Additionally, the computational complexity of integrating causal reasoning into foundation models poses a significant challenge. However, low-rank adaptation methods such as LoRA can be employed to reduce the number of learnable parameters, making causal integration more efficient without compromising performance hu2021lora .
@@ 5 Alternative View
Symbolic AI leverages prior knowledge through logic rules, ontologies, and formal representations, supporting structured reasoning, explainability, and generalization from small data diaz2022explainable . It can also help navigate trade-offs, for example, between accuracy and interoperability, by making the reasoning process more transparent and controllable. While both symbolic AI and causal reasoning rely on prior assumptions, causality uniquely enables reasoning about interventions and counterfactuals. Unlike symbolic systems that model static relationships, causal models capture how changes in one variable affect others - an essential feature in domains like fairness and policy evaluation.
Another perspective suggests that causal properties may emerge spontaneously as models scale and are trained on increasingly diverse data. Recent theoretical work by richens2024 suggests that robust agents implicitly learn causal world models if exposed to sufficiently diverse environments. However, this emergence is neither guaranteed nor controllable. In contrast, explicitly integrating existing scientific causal knowledge into machine learning provides a more reliable, interpretable, and resource-efficient pathway to trustworthy AI.